OpenStudy (anonymous):

simplify: (-sq3sinx - sq3sinx)/2

OpenStudy (anonymous):

\[-\sqrt{3}\sin x - \sqrt{3}\sin x \]

OpenStudy (anonymous):

(−3√sinx−3√sinx)/2